AFWEX Status, 29 November 2000

Held a kick-off science telecon between Tinker, Stillwater, and the CART site at 2pm. The agenda included an overview of the AFWEX objectives, instrument status, and today and tomorrow's plans. At Tinker, Rich Ferrare and John Goldsmith showed preliminary results from the 11/27 ferry flight.

The DC-8 and Proteus flew coordinated flights over the CF with operations between ~1630 and ~2230 CST. The Proteus flew a slow acent followed by a mapping pattern over the CF for three hours and a slow descent. The DC-8 flew a one hour mission for COAST followed by a spiral up centered on the CF, level leg bowties at 5 flight levels between 41 and 25 kft, and a spiral down over the CF. Coincident observations from the CART site were made.

Light cirrus at 25kft at mid-morning with clearing into evening. Severe clear and stable conditions during the flights.

Instrument Status/Comments
CART site:
   CART Raman Lidar Normal operations. Using fixed calibration (determined from WVIOP2000 time frame) throughout AFWEX.
   GSFC Raman Lidar Sub-optimal but Operational. YAG service visit today. Operated from 1800 to 2200 CST.
   MPI DIAL Not operational. Waiting on parts.
   WFF Chilled Mirror Sondes Normal Operations. Launches at 1715 and 2015 CST. 20 units left. Working issue of slow ascent rates.
   Other University of Wisconsin personel made land surface temperature and emissivity measurements in the CART site vicinity throughout the day and during the aircraft overflights.
   LASE Normal operation.
   DLH Normal operation.
   Cryo frost-point/chilled-mirror Normal operation.
   SHIS Normal operation.
   COAST Normal operation.
   AeroSapient Normal operation.
   NAST-I Normal operation.
   NAST-M Operated, but no science data recovered.
   FIRSC Had detectors working and collected first science data.

Dave Tobin, University of Wisconsin